# Recovering orthogonality from quasi-nature of Spectral transformations

### v1

@@ Abstract.
Orthogonal Polynomials and linear functionals, linear spectral transformations, quasi-orthogonal polynomials, recurrence relations, transfer matrix, Stieltjes transform, continued fraction.
@@ 3.1. An alternative representation of $\beta_{n}$
We have started with the sequence of monic orthogonal polynomials $\{P_{n}(x)\}_{n=0}^{\infty}$ with respect to the quasi-definite linear functional $\mathcal{L}$. Linear spectral transformations of $\mathcal{L}$ yield several sequences of orthogonal polynomials. From the Geronimus polynomials, we introduce the concept of quasi-Geronimus polynomial of order one. In this process, the orthogonality condition for the quasi-Geronimus polynomial of order one is relaxed. However, the previous result indicates that orthogonality can still be achieved with a suitable sequence of the constants $\beta_{n}$ in the definition of monic quasi-Geronimus polynomials.
@@ 5.1. The Geronimus Case
The monic quasi-Geronimus polynomial of order one is given by
$\mathcal{G}^{Q}_{n+1}(x;0)=\mathcal{G}_{n+1}(x;0)+\beta_{n+1}\mathcal{G}_{n}(x;0).$
@@
$\beta_{n}=0.5$, $\alpha=0.9$ $\beta_{n}=1$, $\alpha=1.5$ $\beta_{n}=-1.5$, $\alpha=1$ $\beta_{n}=-2$, $\alpha=0.5$
0.20772 0.39721 -1.22375 -3.62415
1.19735 1.61551 0.35502 0.13763
3.08547 3.74594 1.76306 1.23604
6.04014 6.99107 4.23274 3.46801
10.4397 11.83655 8.00906 7.05233
17.4297 20.41380 13.8639 12.7301
@@
We see that for $\beta_{n}=0.5$ and $\alpha=0.9$, all zeros of $\mathcal{G}^{Q}_{6}(x;0)$ are within the support of the Laguerre weight. Similarly, for $\beta_{n}=1$ and $\alpha=1.5$, the zeros also lie within the support. However, when $\beta_{n}=-1.5$ and $\alpha=1$, exactly one zero of $\mathcal{G}^{Q}_{6}(x;0)$ is outside the support, as it is shown in Table 1. The same holds true for $\beta_{n}=-2$ and $\alpha=0.5$.
@@
By choosing $\beta_{1}=\alpha$ and $\beta_{2}=\alpha+1$, we recursively obtain $\beta_{n}=n+\alpha-1$. Thus, by Proposition 1, the monic quasi-Geronimus polynomial of order one, denoted as $\mathcal{G}^{Q}_{n+1}(x;0)$, given by
$\mathcal{G}^{Q}_{n+1}(x;0)=\mathcal{G}_{n+1}(x;0)+(n+\alpha)\mathcal{G}_{n}(x;0),$
@@
As observed, the sequence of quasi-Geronimus Laguerre polynomial of order one is orthogonal when $\beta_{n}=n+\alpha$. Consequently, the zeros of $\mathcal{G}^{Q}_{n+1}(x;0)$ are within the interval $(0,\infty)$. In Table 2, we illustrate that, for $\alpha=1$ and $\alpha=0.5$, the zeros of $\mathcal{G}^{Q}_{5}(x;0)$ lie within the interval $(0,\infty)$. Additionally, Figure 1 and Table 3 show that for $\alpha=1$, the zeros of $\mathcal{G}^{Q}_{5}(x;0)$ and $\mathcal{G}^{Q}_{6}(x;0)$ interlace.
@@
$\alpha=1$, $n=4$, $\beta_{n+1}=n+\alpha$ $\alpha=0.5$, $n=4$, $\beta_{n+1}=n+\alpha$
0.31192 0.140056
1.68706 1.28981
4.3774 3.77609
9.03713 8.22922
34.5865 31.5648
@@ 6. Concluding remarks
We derive the closed form of $\beta_{n}$ satisfying (3.9), necessary for the orthogonality of quasi-Geronimus Laguerre polynomials of order one. Specifically, the recurrence parameters and the three-term recurrence relation satisfied by the quasi-Geronimus Laguerre polynomial of order one are obtained. Moreover, it would be of interest to determine, if possible, the explicit form of $\alpha_{n}$ satisfying (4.9) and to recover the recurrence coefficients that ensure the existence of an orthogonality measure for the quasi-Uvarov Laguerre polynomial of order one.

### v2

@@ Abstract.
Orthogonal Polynomials and linear functionals; linear spectral transformations; quasi-orthogonal polynomials; recurrence relations; transfer matrix; Stieltjes transform; continued fraction.
@@ 3.1. An alternative representation of $\beta_{n}$
We have started with the sequence of monic orthogonal polynomials $\{P_{n}(x)\}_{n=0}^{\infty}$ with respect to the quasi-definite linear functional $\mathcal{L}$. Linear spectral transformations of $\mathcal{L}$ yield several sequences of orthogonal polynomials. From the Geronimus polynomials, we introduce the concept of quasi-Geronimus polynomial of order one. In this process, the orthogonality condition for the quasi-Geronimus polynomial of order one is relaxed. However, the previous result indicates that orthogonality can still be achieved with a suitable sequence of the constants $\beta_{n}$ in the definition the monic quasi-Geronimus polynomials.
@@ 5.1. The Geronimus Case
The monic quasi-Geronimus Laguerre polynomial of order one is given by
$\mathcal{G}^{Q}_{n+1}(x;0)=\mathcal{G}_{n+1}(x;0)+\beta_{n+1}\mathcal{G}_{n}(x;0)=\mathcal{L}^{(\alpha-1)}_{n+1}(x)+\beta_{n+1}\mathcal{L}^{(\alpha-1)}_{n}(x).$ (5.9)
@@
$\beta_{n}=0.5$, $\alpha=0.9$ $\beta_{n}=1$, $\alpha=1.5$ $\beta_{n}=7$, $\alpha=1$ $\beta_{n}=6$, $\alpha=0.1$
0.193294 0.355981 -0.248125 -0.0584409
1.11293 1.44484 0.475247 0.108916
2.86119 3.3362 1.96233 1.31668
5.57689 6.17578 4.45828 3.53458
9.5578 10.2685 8.24579 7.05567
15.5979 16.4187 14.1065 12.6426
@@
We see that for $\beta_{n}=0.5$ and $\alpha=0.9$, all zeros of $\mathcal{G}^{Q}_{6}(x;0)$ are within the support of the Laguerre weight. Similarly, for $\beta_{n}=1$ and $\alpha=1.5$, the zeros also lie within the support. However, when $\beta_{n}=7$ and $\alpha=1$, exactly one zero of $\mathcal{G}^{Q}_{6}(x;0)$ is outside the support, as it is shown in Table 1. The same holds true for $\beta_{n}=6$ and $\alpha=0.1$.
@@
By choosing $\beta_{1}=\alpha$ and $\beta_{2}=\alpha+1$, we recursively obtain $\beta_{n}=n+\alpha-1$. Thus, by Proposition 1, the monic quasi-Geronimus Laguerre polynomial of order one, denoted as $\mathcal{G}^{Q}_{n+1}(x;0)$, given by
$\mathcal{G}^{Q}_{n+1}(x;0)=\mathcal{G}_{n+1}(x;0)+(n+\alpha)\mathcal{G}_{n}(x;0)=\mathcal{L}^{(\alpha-1)}_{n+1}(x)+(n+\alpha)\mathcal{L}^{(\alpha-1)}_{n}(x),$ (5.11)
@@
As observed, the sequence of quasi-Geronimus Laguerre polynomial of order one is orthogonal when $\beta_{n+1}=n+\alpha$. Consequently, the zeros of $\mathcal{G}^{Q}_{n+1}(x;0)$ are within the interval $(0,\infty)$. In Table 2, we illustrate that, for $\alpha=0.1$ and $\alpha=0.5$, the zeros of $\mathcal{G}^{Q}_{5}(x;0)$ lie within the interval $(0,\infty)$, with one zero positioned exactly on the boundary of the support. Additionally, Figure 1 and Table 3 show that for $\alpha=1$, the zeros of $\mathcal{G}^{Q}_{5}(x;0)$ and $\mathcal{G}^{Q}_{6}(x;0)$ interlace within the interval $(0,\infty)$.
@@
$\alpha=0.1$, $n=4$, $\beta_{n+1}=n+\alpha$ $\alpha=0.5$, $n=4$, $\beta_{n+1}=n+\alpha$
0 0
0.36103 0.523526
1.8276 2.15665
4.65741 5.13739
9.55395 10.1824
@@
For further analysis, we consider the monic case of the quasi-Geronimus polynomial of order one $\mathcal{G}_{n}^{Q}(x;a)$ given in (3.1), where $b_{n}=1$. Hence (3.1) can be rewritten as:
$\mathcal{G}_{n}^{Q}(x;a)=\mathcal{G}_{n}(x;a)+k_{n}\mathcal{G}_{n-1}(x;a).$ (5.12)
Note that $\beta_{n}$ is replaced by $k_{n}$, because specific condition on $\beta_{n}$ given by (3.3) turns the quasi-Geronimus polynomial of order one $\mathcal{G}_{n}^{Q}(x;a)$ to quasi-Geronimus orthogonal polynomial of order one, which we denote as $\widetilde{\mathcal{G}}_{n}^{Q}(x;a)$. Note that the polynomial $\widetilde{\mathcal{G}}_{n}^{Q}(x;a)$ satisfies Proposition 1, with condition on $\beta_{n}$ given by (3.3). Consequently, the coefficient $\beta_{n+1}$ involved in (5.9) should be replaced by $k_{n+1}$, resulting in the modified expression:
$\mathcal{G}^{Q}_{n+1}(x;0)=\mathcal{G}_{n+1}(x;0)+k_{n+1}\mathcal{G}_{n}(x;0)=\mathcal{L}^{(\alpha-1)}_{n+1}(x)+k_{n+1}\mathcal{L}^{(\alpha-1)}_{n}(x).$ (5.13)
Similarly, we can rewrite quasi-Geronimus orthogonal Laguerre polynomial of order one (5.11) as:
$\widetilde{\mathcal{G}}^{Q}_{n+1}(x;0)=\mathcal{G}_{n+1}(x;0)+\beta_{n+1}\mathcal{G}_{n}(x;0)=\mathcal{L}^{(\alpha-1)}_{n+1}(x)+(n+\alpha)\mathcal{L}^{(\alpha-1)}_{n}(x).$ (5.14)
We know that at most one zero of a quasi-orthogonal polynomial of order one lies outside the interval of orthogonality [9, Theorem 5.2]. Specifically, Table 1 illustrates that either exactly one zero of the $\mathcal{G}^{Q}_{n}(x;0)$ defined in (5.13) can be negative, or all zeros of the $\mathcal{G}^{Q}_{n}(x;0)$ defined in (5.13) can be positive. However, it is still unclear what conditions are necessary for the coefficient $k_{n}$, as indicated in (5.12), to result in exactly one zero lying outside the interval of orthogonality. Similarly, it is unknown under what circumstances all zeros lie within this interval without transforming the quasi-Geronimus polynomial of order one $\mathcal{G}^{Q}_{n}(x;a)$ into an orthogonal system.
To address this question, we graphically analyze the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ defined in (5.13). We provide two cases to elucidate this matter further:
$\alpha=2.5$, $n=5$
Zeros of $\mathcal{G}^{Q}_{n}(x;0)$, $k_{5}=25$ Zeros of $\widetilde{\mathcal{G}}^{Q}_{n}(x;0)$, $\beta_{n}=n+\alpha-1$
-15.4435 0
1.02900 1.48624
3.13601 3.83768
6.60152 7.48206
12.1768 13.194
Table 5. $\mathcal{G}^{Q}_{n}(x;0)$(5.13) and $\widetilde{\mathcal{G}}^{Q}_{n}(x;0)$ (5.14).
$\alpha=3$, $n=6$
Zeros of $\mathcal{G}^{Q}_{n}(x;0)$, $k_{6}=10$ Zeros of $\widetilde{\mathcal{G}}^{Q}_{n}(x;0)$, $\beta_{n}=n+\alpha-1$
-0.966814 0
1.30868 1.49055
3.37753 3.58133
6.41475 6.627
10.728 10.9444
17.1378 17.3567
Table 6. $\mathcal{G}^{Q}_{n}(x;0)$(5.13) and $\widetilde{\mathcal{G}}^{Q}_{n}(x;0)$ (5.14).
(a) Graph of $\mathcal{L}_{n}^{(\alpha)}(x)$ (red), $\mathcal{L}_{n}^{(\alpha-1)}(x)$ (orange), $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) and $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (yellow). $n=5,\alpha=2.5,k_{5}=25$.
(b) Graph of $\mathcal{L}_{n}^{(\alpha)}(x)$ (red), $\mathcal{L}_{n}^{(\alpha-1)}(x)$ (orange), $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) and $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (yellow). $n=6,\alpha=3,k_{6}=10$.
Figure 3. $\mathcal{G}_{n}^{Q}(x;0)$ (5.13) and $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (5.14). For $k_{n}>n+\alpha-1$.
Case 1: when $\bf{k_{n}>n+\alpha-1}$. Figure 3 illustrates that when the coefficient $k_{n}$ given in (5.13) exceeds $n+\alpha-1$, for any fixed degree, the growth of the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) defined in (5.13) surpasses the growth of the polynomial $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (yellow) defined in (5.14) beyond the last intersection point of the corresponding polynomials. It is worth mentioning that a negative zero of the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ arises only when $k_{n}$ exceeds $n+\alpha-1$. If $k_{n}$ is significantly larger than $n+\alpha-1$, this negative zero moves away from the first zero (i.e., $x=0$) of the polynomial $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ at a faster rate. Whereas, if $k_{n}$ is greater but closer to $n+\alpha-1$, the negative zero remain in the vicinity of the origin. This phenomenon is illustrated in Table 5 and Table 6 for $n=5,6$. However, the scenario changes entirely when we examine the growth of these two polynomials before reaching the first intersection point. For odd-degree polynomials, the graph of $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (yellow) always remains below the graph of $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) prior to the first intersection point of the polynomials, as depicted in Figure 3(a). On the other hand, for even degrees, the graph of $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (yellow) grows faster than the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) before reaching the first intersection point of the corresponding polynomials, as demonstrated in Figure 3(b).
$\alpha=4.5$, $n=5$
Zeros of $\mathcal{G}^{Q}_{n}(x;0)$, $k_{5}=3$ Zeros of $\widetilde{\mathcal{G}}^{Q}_{n}(x;0)$, $\beta_{n}=n+\alpha-1$
1.54164 0
3.5791 2.61181
6.54479 5.56339
10.7668 9.76859
17.0676 16.0562
Table 7. $\mathcal{G}^{Q}_{n}(x;0)$(5.13) and $\widetilde{\mathcal{G}}^{Q}_{n}(x;0)$ (5.14).
$\alpha=3$, $n=4$
Zeros of $\mathcal{G}^{Q}_{n}(x;0)$, $k_{4}=2$ Zeros of $\widetilde{\mathcal{G}}^{Q}_{n}(x;0)$, $\beta_{n}=n+\alpha-1$
1.07821 0
3.05525 2.14122
6.29734 5.31552
11.5692 10.5433
Table 8. $\mathcal{G}^{Q}_{n}(x;0)$(5.13) and $\widetilde{\mathcal{G}}^{Q}_{n}(x;0)$ (5.14).
(a) Graph of $\mathcal{L}_{n}^{(\alpha)}(x)$ (red), $\mathcal{L}_{n}^{(\alpha-1)}(x)$ (orange), $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) and $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (yellow). $n=5,\alpha=4.5,k_{5}=3$.
(b) Graph of $\mathcal{L}_{n}^{(\alpha)}(x)$ (red), $\mathcal{L}_{n}^{(\alpha-1)}(x)$ (orange), $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) and $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (yellow). $n=4,\alpha=3,k_{4}=2$.
Figure 4. $\mathcal{G}_{n}^{Q}(x;0)$ (5.13) and $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (5.14). For $k_{n}<n+\alpha-1$.
Case 2: when $\bf{k_{n}<n+\alpha-1}$. The situation differs significantly compared to Case 1. It is evident from Table 7 and Table 8 that no negative zero of the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ defined in (5.13) occurs when $\beta_{n}<n+\alpha-1$. In other words, all the zeros of the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ lie within the interval $(0,\infty)$. For any fixed degree, Table 7 and Table 8 illustrate that the first zero of the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ does not fluctuate significantly. As expected from Case 1, the polynomial $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (yellow) tends to dominate the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) after the last intersection point when $k_{n}<n+\alpha-1$. For odd degrees, the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) defined in (5.13) decays faster than the polynomial $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$ (yellow) defined in (5.14) before reaching the first intersection point of the polynomials, as illustrated in Figure 4(a). On the other hand, for even degrees, the growth of the polynomial $\mathcal{G}_{n}^{Q}(x;0)$ (magenta) exceeds that of $\widetilde{\mathcal{G}}_{n}^{Q}(x;0)$(yellow) before reaching the first intersection point of the polynomials, as demonstrated in Figure 4(b).
@@ 6. Concluding remarks
We derived the closed form of $\beta_{n}$ satisfying (3.3), necessary for the orthogonality of the quasi-Geronimus Laguerre polynomials of order one $\mathcal{G}_{n}^{Q}(x;a)$. Specifically, the recurrence parameters and the three-term recurrence relation satisfied by the quasi-Geronimus Laguerre polynomial of order one are obtained. Moreover, it would be of interest to determine, if possible, the explicit form of $\alpha_{n}$ satisfying (4.5) and to recover the recurrence coefficients that ensure the existence of an orthogonality measure for the quasi-Uvarov Laguerre polynomial of order one. Finally, drawing from the numerical experiments conducted in Section 5, we conclude this manuscript by summarizing the following observations::
Observation 1.
Let $\mathcal{G}_{n}^{Q}(x;a)$ be the polynomial of degree $n$ as defined in (5.12), with a non-zero unknown parameter $k_{n}$. Let $\widetilde{\mathcal{G}}_{n}^{Q}(x;a)$ represent the degree $n$ monic quasi-Geronimus orthogonal polynomial of order one, with the known parameter $\beta_{n}$ provided in Proposition 1. If $\mathcal{G}_{n}^{Q}(x;a)$ and $\widetilde{\mathcal{G}}_{n}^{Q}(x;a)$ intersect exactly at $m$ points, which are ordered as
$x_{1}<x_{2}<...<x_{m},$
then for any fixed degree $n$ and for any $k_{n}>\beta_{n}$, we have
$\widetilde{\mathcal{G}}_{n}^{Q}(x;a)<\mathcal{G}_{n}^{Q}(x;a),$
whenever $x>x_{m}$. Moreover, for $n=2k$,
$\widetilde{\mathcal{G}}_{2k}^{Q}(x;a)>\mathcal{G}_{2k}^{Q}(x;a),$
and for $n=2k-1$,
$\widetilde{\mathcal{G}}_{2k}^{Q}(x;a)<\mathcal{G}_{2k}^{Q}(x;a),$
whenever $x<x_{1}$.
Observation 2.
Let $\mathcal{G}_{n}^{Q}(x;a)$ denote the monic quasi-Geronimus polynomial of order one, as defined in (5.12), with a free non-zero parameter $k_{n}$. The coefficient $\beta_{n}$ is given by Proposition 1. Then the following statements hold:
  1. (1)
    If $k_{n}>\beta_{n}$, then exactly one zero of $\mathcal{G}_{n}^{Q}(x;a)$ lies outside the interval of orthogonality.
  2. (2)
    If $k_{n}<\beta_{n}$, then all the zeros of $\mathcal{G}_{n}^{Q}(x;a)$ lies inside the interval of orthogonality.